OpenStudy (anonymous):

A community group sells 2,000 tickets for its raffle. The grand prize is a car. Neil and his 9 friends buy 10 tickets each. When the winning ticket number is announced, it is found to belong to Neil's group. Given this information, what is the probability that the ticket belongs to Neil?

OpenStudy (anonymous):

90/2000

OpenStudy (anonymous):

90/2000??????

OpenStudy (dangerousjesse):

What's \(9\times 10\)?

OpenStudy (anonymous):

justify @DangerousJesse m too confused.

OpenStudy (anonymous):

10/90

OpenStudy (anonymous):

got it..ya..

OpenStudy (texaschic101):

neil and his 9 friends buy 10 tickets each.....thats 10 x 10 = 100 tickets in neils group but for the probability for it to be Neils...10/100 = 1/10

OpenStudy (anonymous):

@Probhat not only are your answers both wrong, you also should not just post an answer without explanation.

OpenStudy (anonymous):

simple 1/2000

OpenStudy (anonymous):

reading and understanding the question is important if it is not clear what it is asking, don't try to do it

OpenStudy (anonymous):

and no. @parikpelletpro it is not that

OpenStudy (anonymous):

ur right

OpenStudy (anonymous):

go with @texaschic101

OpenStudy (anonymous):

neil is one of ten people go with \(\frac{1}{10}\)

OpenStudy (anonymous):

10/100 is correct

OpenStudy (anonymous):

texaschic101 verified proof of answer: http://prntscr.com/7cu9i2